Consider the reaction, S2O82-   +    3 I-  à    2 SO42-  +  I3-. Data obtained in measuring rate of formation of I3- are listed in the table.

Experiment

[S2O82-], M

[I], M

Initial rate, Ms-1

1

0.035

0.055

1.5 x 10-5

2

0.070

0.055

3.0 x 10-5

3

0.070

0.110

6.0 x 10-5

  1. Determine the order of reaction with respect to S2O82-.
  1. What is the order of reaction with respect to I-?
  1. Give the overall order of reaction.
  1. Write the rate law for the reaction.
  1. Determine the value for the rate constant, k.
  1. What is the initial rate of reaction if [S2O82-] = 0.030M and [I-] = 0.050 M?
  1. When the rate of formation of I3-is 3.0 x 10-5 M/s, what is the rate of disappearance of I-?
Answer #1

Solution :-

Suppose, the order with respect to S2O82- is p and with respect to I- is q. The rate law can thus be written as: Rate = k [S​​​​​​​​​​​​​​2​​​​​​​O​​​​​​​8]p [I-]q

Substituting the values of rate, [S​​​​​​​2​​​​​​​O​​​​​​​8] and [I-] for various sets of experiments, we have

1.5 × 10-5 = k [0.035]p [0.055]q ............….(i)

​​​​​​3.0 × 10-5 = k [0.070]p [0.055]q ................(ii)

6.0 × 10-5 = k [0.070]p [0.110]q ...............(iii)

Dividing eq. (ii) by eq. (i) , we get,

(3.0 × 10-5) / (1.5 × 10-5) = (0.070)p / (0.035)p

2 = 2p

p = 1

Dividing eq. (iii) by eq. (ii), we get,

(6.0 × 10-5) / (3.0 × 10-5) = (0.110)q / (0.055)q

2 = 2q

q = 1

1) Thus, the order of reaction with respect to S2O82- is = 1

2) The order of reaction with respect to l- is = 1

3) Overall rate of reaction = (p + q) = (1 + 1) = 2

4) Rate Law ,

Rate = k [S​​​​​​​2​​​​​​​O​​​​​​​8​​​​​​​​2-​​​​​​]1 [I-]1

5) Putting values of order of reactants in equation 1, we get,

1.5 × 10-5 = k [0.035]1 [0.055]1

or, k (Rate constant) = (1.5 × 10-5) / (0.035 × 0.055) = 7.8 × 10-3  M-1 s-1

6) If [S2O82-​​​​​​] = 0.030 M, and [I​​​​​​​-] = 0.050 M

k = 7.8 × 10-3  M-1 s-1

Rate =?

Rate = (7.8 × 10-3) × (0.030)× (0.050)

Rate = 1.17 × 10-5 M s-1

7) Rate of formation of I3- = 1/3 × Rate of disappearance of I-

or , Rate of disappearance of I- = 3 × Rate of formation of I3-

or, Rate of disappearance of I​​​​​​​- = 3 × (3.0 × 10-5) = 9.0 × 10-5 M/s

-----------------------------------------------------------------------------